# Inferring fine-grained migration patterns across the United States

### v1

@@ 4 Case study: migration in response to California wildfires
We demonstrate the value of the MIGRATE estimates by using them to study migration in response to wildfires in California. Natural disasters, including wildfires, are known drivers of human mobility [3, 26]. There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area. In many US states, including California, wildfire risk increased from 2010-2019 [27], and is expected to further increase due to climate change [28]. Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters [29].
@@
MIGRATE estimates reveal dramatic levels of out-migration for CBGs which lie within the fire perimeters (Figure 4a) in the year following the fires, often exceeding 50%. [^4] The fact that the out-migration rate is not 100% is likely due to a number of factors, including the often-preferred option of rebuilding and returning [37, 38]. Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [29]. In contrast, CBGs which do not lie within the fire perimeters experience much lower out-migration rates. We systematically quantify these differences in Figure 4b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole. The out-migration rate in the year following the Camp Fire for CBGs within the fire perimeter is 46%, at least 3.1$\times$ that of less-affected CBG groups; the out-migration rate in the year following the Tubbs Fire for CBGs within the perimeter is 37%, at least 2.8$\times$ that of less-affected CBG groups. Our estimates of out-migration following the Camp Fire are similar to those in prior work [26].
@@
Collectively, these analyses showcase how MIGRATE estimates reveal substantial, policy-relevant responses to wildfires which traditional data sources conceal. Indeed, the MIGRATE estimates enable further analyses beyond those conducted here: for example, they can illuminate inequality in responses to wildfires by connecting out-migration patterns with CBG-level demographics.
@@ 5 Discussion
We produce MIGRATE, a dataset of spatially and temporally fine-grained migration matrices capturing annual flows between pairs of CBGs from 2010-2019. Our estimates are approximately 4,600 times more spatially granular than publicly available county-to-county 5-year migration data, and 18 million times more spatially granular than state-to-state 1-year migration data. We produce these estimates by developing an IPF-based algorithm to harmonize granular but biased proprietary migration data with more reliable but less granular Census datasets.
We show that MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to the raw proprietary data. We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix F. Our estimates are available at migrate.tech.cornell.edu.
@@
While we validate MIGRATE estimates comprehensively against external data sources, practitioners making use of the data should be mindful of two limitations. First, while we use multiple ground-truth Census data sources to reduce the biases we document in the raw Infutor data, uncorrected biases likely remain. For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows. Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets which are not included in producing our estimates), these validations have imperfections. As we further describe in Appendix B.2, Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other. Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model [48, 49]. Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
@@
We hope these improvements will facilitate a wide range of further migration-related analyses. As our wildfire case study illustrates, MIGRATE reveals patterns which remain invisible in publicly available migration data. The fine spatial scale of the data also enables analyses of inequality in migration, by linking migration patterns to CBG-level demographics. Our work also highlights methodological directions for improving estimation of fine-grained migration patterns, including providing principled margins of error on estimates and incorporating additional data sets to further improve estimation and validation. More broadly, we hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
Appendix A Processing Infutor Data
@@
Infutor provides data in tabular form. Each row in the data provides data for one individual, which includes a list of known addresses, along with the date when the individual is first observed at the address (which we refer to the effective date below), and the first and last date that the individual is observed in the data (which we refer to as the listed start date and the listed end date, respectively). [^5] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information). We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known biases [20]. Dates are listed as month and year. For example, the data for one individual might consist of two addresses: “1 Main Street, Everytown, USA, 10000 (January 2010); 2 Cornelia Street, New York, New York, USA (December 2017)” followed by an initial date of January 2008 and an end date of December 2017. We describe the steps taken to (1) identify and clean Infutor data records within the scope of MIGRATE, (2) process the Infutor data into matrices documenting yearly flows between address pairs, and (3) map these address-level matrices to the CBG-level matrices $E^{(t)}$ mentioned in the main text. Table 2 summarizes the raw and processed data.
A.1 Cleaning address histories
@@
A.2 Processing address histories into annual address-address migration matrices
@@
A.3 Processing address-address migration matrices into CBG-CBG migration matrices
Figure 5: Flowchart detailing the process of mapping addresses to CBGs. We start with all Infutor addresses and classify them into five address types. We remove addresses which lie within US territories but not US states (around .26% of the addresses). Incomplete addresses—those that do not have a precise street address line—as well as PO boxes and rural routes are mapped probabilistically to CBGs according to their ZIP code. The vast majority (90.4%) of addresses are clean, complete street addresses which are sent to the Census geocoder, which is able to map 81.26% of these addresses to a Census block group. If the Census geocoder fails to provide a match, we reattempt matching using ESRI’s ArcGIS geocoder; this achieves a lower match rate of 58.24%, in part because the sample it is applied to is more difficult to parse. In total, we are able to map 99.21% of all addresses in the original Infutor data: 83.33% to a precise latitude and longitude, and 15.89% to a ZIP code.
We transform the address-level migration matrices $A^{(t)}$ into CBG-level matrices $E^{(t)}$ by mapping each address either to a precise CBG where possible, or to a distribution over CBGs when only a ZIP code is available. We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency. Figure 5 details this process.
Our address mapping pipeline is divided in two parts: mapping addresses via geocoding, and mapping addresses with a ZIP code-to-CBG crosswalk. We initially attempt to use one of two state-of-the art geocoders to map addresses to latitudes and longitudes: the publicly-available Census Bureau geocoder [23] and ESRI’s ArcGIS geocoder [22]. We first attempt to geocode an address via the Census geocoder; if this algorithm does not find a match, we submit the address to ArcGIS. If neither algorithm finds a match, we map the address to Census Tracts intersecting its ZIP code, weighting probabilities based on the share of residential ZIP code addresses represented by the population of each Census Tract. Addresses representing rural routes and postal boxes are directly mapped via ZIP code, as well as incomplete addresses. We use the HUD ZIP-to-tract crosswalk, with the date closest available to the last seen date of an address to account for the fact that ZIP code definitions may vary and are not nested within Census geographies [50]. We then distribute the probability to each Census block group within the tracts, again weighting by their relative population shares. Success rates for each of these steps are high and detailed in Figure 5.
@@ A.3 Processing address-address migration matrices into CBG-CBG migration matrices
Table 2 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline. The number of individuals with active records in the data represents the number of individuals with an interval of activity containing each year at the start of the data processing; the number of CBG moves corresponds to the sum of off-diagonal entries of each matrix $E^{(t+1)}$ and measures the “effective size” of each year’s data: how many movers we actually account for at the beginning of the estimation procedure.
Year (1) Active records (2) US Population (3) Effective Dates (4) Processed CBG Moves
2010 269,228,776 309,327,143 48,225,002 15,180,982
2011 260,064,921 311,583,481 28,196,562 11,793,806
2012 259,451,204 313,877,662 32,009,822 12,544,199
2013 252,933,953 316,059,947 27,386,564 13,829,056
2014 256,453,630 318,386,329 37,386,684 13,071,070
2015 241,295,957 320,738,994 30,777,204 14,022,249
2016 242,298,262 323,071,755 38,085,337 13,605,804
2017 193,479,001 325,122,128 34,599,168 12,684,184
2018 177,629,604 326,838,199 33,183,982 13,505,890
Table 2: Yearly Breakdown of Data Summaries. Population and addresses accounted by the Infutor and Census datasets during the analysis period (2010-2019). (1) Individuals have active records in a given year if the year falls within their interval of activity in the dataset. The number of active individuals decreases in recent years, as Infutor records tend to appear when a move is reported. Comparison of these values to (2) the Census population shows that Infutor under-counts the population at every given year. (3) The number of effective dates per year is naturally reduced as we process the data into (4) CBG-to-CBG moves: an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
Appendix B Processing Census Data
The U.S. Census Bureau runs multiple programs with the aim of counting and estimating demographic and socioeconomic population data. Each of these programs may release data at different spatiotemporal granularity and, even if geographies or timespans agree, the data might lack internal consistency. We use a series of steps to clean, select, and process Census datasets. We describe our procedures for reconciling Census geographies (§B.1), selecting which Census datasets to use as constraints (§B.2), and how we process the Census data to deal consistently with births, deaths, emigration, and immigration (§B.3).
B.1 Reconciling Census geographies
@@
B.2 Selecting Census datasets
@@
Table 3: ACS Coefficients of Variation. Average coefficient of variation (in %) of the non-zero estimates in each dataset. CVs were computed as the ratio of the standard error (obtained by dividing the margin of error by $1.645$) by the estimate [54]. We did not use flows released in 2010 (as they would report moves from 2009).
@@
We would like to avoid matching to datasets that were too noisily estimated. To decide which ACS datasets to use in our harmonization process, we examine their level of sampling error. We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value. Table 3 reports the mean coefficient of variation in ACS estimates. In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%). We opt not to match to county-county flows due to the imprecision of raw flow estimates, although we do use them as a validation. In-migration rates have much tighter CVs than the corresponding flows (mean around 15% and 5% for counties and states, respectively), and we also use those as validations.
@@
B.3 Accounting for natural population increase and international migration
@@
Appendix C Harmonizing Infutor and Census Data
@@
C.1 Harmonizing with CBG population data
@@
C.2 Harmonizing with yearly state non-movers data
@@
C.3 Harmonizing with yearly state-to-state flow data
@@
C.4 Harmonizing with yearly county population data
@@
Appendix D Validations
Here we provide additional details on MIGRATE validations. §D.1 provides additional validation metrics that complement Figure 2; §D.2 presents additional details on our held-out data validations; §D.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §D.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
D.1 Additional validation metrics
Figure 6: Detailed flow validations of MIGRATE. To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
Figure 6 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We report these metrics at both the state and county level; state-to-state flows were used in the estimation procedure, county-to-county flows were held out. For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
D.2 Validating the estimation procedure using held-out data

### v2

@@ 2.4 Analysis of local migration patterns
In addition to using MIGRATE to analyze national migration trends, we use it to study local migration patterns: specifically, migration in response to wildfires in California. Natural disasters, including wildfires, are known drivers of human mobility [3, 29]. There were over 3,000 fire events in California from 2010-2019, which cumulatively affected nearly 27,504 square kilometers — approximately 6.5% of the California land area. In many US states, including California, wildfire risk increased from 2010-2019 [30], and is expected to further increase due to climate change [31]. Researchers rely on a variety of data sources, such as administrative records and building codes, to study vacancy and movement following these disasters [32].
@@
MIGRATE estimates reveal dramatic levels of out-migration for CBGs which lie within the fire perimeters (Figure 5a) in the year following the fires, often exceeding 50%. [^6] The fact that the out-migration rate is not 100% is likely due to a number of factors, including the often-preferred option of rebuilding and returning [40, 41]. Looking at administrative address data from mail delivery after Camp Fire, for example, researchers found a much more significant uptick in short-term vacancy than in long-term vacancy [32]. In contrast, CBGs that do not lie within the fire perimeters experience much lower out-migration rates. We systematically quantify these differences in Figure 5b, which compares CBGs within the fire perimeter to three sets of less-affected CBGs outside the perimeter: other CBGs neighboring the perimeter; other CBGs within the affected counties; and others within California as a whole. The out-migration rate in the year following the Camp Fire for CBGs within the fire perimeter is 46%, at least 3.1$\times$ that of less-affected CBG groups; the out-migration rate in the year following the Tubbs Fire for CBGs within the perimeter is 37%, at least 2.8$\times$ that of less-affected CBG groups. Our estimates of out-migration following the Camp Fire are similar to those in prior work [29].
@@
In Appendices C and D, we provide two additional analyses of local migration patterns that are impossible using only county-level data: we analyze socioeconomic variation in New York City out-mover destinations, and migration patterns for residents who live in public housing provided by the New York City Housing Authority as part of the city’s affordable housing policy. Collectively, these analyses showcase how MIGRATE estimates reveal important, policy-relevant local migration patterns that traditional data sources conceal.
@@ 3 Discussion
We produce MIGRATE, a dataset of spatially and temporally fine-grained migration matrices capturing annual flows between pairs of CBGs from 2010-2019. Our estimates are approximately 4,600 times more spatially granular than publicly available county-to-county 5-year migration data, and 18 million times more spatially granular than state-to-state 1-year migration data. MIGRATE estimates correlate highly with external Census ground-truth datasets and reduce error and demographic bias relative to raw proprietary data. We discuss the measures taken to protect the privacy of all individuals in the dataset in Appendix E. Our estimates are available at migrate.tech.cornell.edu.
We use MIGRATE to analyze both local and national patterns of migration. We show that MIGRATE can reveal important local migration patterns, including dramatic increases in out-migration in response to California wildfires, that are invisible in county-level data. Nationally, we analyze demographic and temporal variation in migration patterns. We find that people tend to move to CBGs of the same type across dimensions of race, income, and rural/urban status — for example, movers from plurality Black, Asian, and Hispanic CBGs are 15.5$\times$, 14.6$\times$, and 4.3$\times$ likelier to move to CBGs with the same plurality race group, consistent with prior work documenting the persistence of racial segregation [51, 52, 53]. We document demographic and temporal variation in moving distance: movers from plurality white, rural, and higher-income CBGs are likelier to travel long distances, and moving distance increases over the decade we study, consistent with prior work [15].
An important finding from our national analysis is racial and temporal variation in upward mobility. Movers are likelier to move to top-income-quartile CBGs than bottom-income-quartile CBGs, a trend that becomes more pronounced over the decade we study. We find large racial disparities in the likelihood of moving to higher-income CBGs, or top-income-quartile CBGs, even when controlling for income of origin CBG: movers from plurality Asian CBGs are likelier to move to a higher-income CBG, and movers from plurality Black CBGs are less likely. These findings are consistent with prior work documenting racial disparities in neighborhood attainment [54, 55, 56, 57] and generalize these findings to larger and more recent samples and additional racial groups. Overall, we document demographic and temporal variation in national migration trends using a recent, large-scale, and highly granular dataset.
@@
While we validate MIGRATE estimates comprehensively against external data sources, practitioners making use of the data should be mindful of two limitations. First, while we use multiple ground-truth Census data sources to reduce the biases we document in the raw Infutor data, uncorrected biases likely remain. For example, while we correct biases in CBG populations, we cannot correct sub-CBG-level population biases, or biases in CBG-CBG flows. Second, while we show that MIGRATE estimates are highly correlated with external Census datasets (including held-out datasets that are not used to produce our estimates), these validations have imperfections. As we further describe in the Methods Section (§M.2), Census datasets themselves have biases; can possess significant margins of error, particularly at fine spatial scales; and are not perfectly consistent with each other. Another challenge in validation is the lack of a ground-truth CBG-CBG migration matrix against which to validate; publicly available Census datasets do not afford this level of granularity, and non-Census data sources (e.g., voter files) have significant biases and measure substantively different populations than the one we seek to model [58, 59]. Overall, MIGRATE should be viewed as a new data source which like all data possesses limitations but which nonetheless represents a significant improvement on widely used publicly available data sources (due to its granularity) or proprietary data sources (due to reductions in error and bias).
@@
We hope these improvements will facilitate a wide range of further migration-related analyses. As our analyses illustrate, MIGRATE reveals patterns which remain invisible in publicly available, county-level migration data. We hope MIGRATE will facilitate more precise study of many migration-related phenomena across the social, health, and environmental sciences.
Methods
M.1 Processing Infutor Data
@@ Methods
Infutor provides data in tabular form. Each row in the data provides data for one individual, which includes a list of known addresses, along with the date when the individual is first observed at the address (which we refer to the effective date below), and the first and last date that the individual is observed in the data (which we refer to as the listed start date and the listed end date, respectively). [^7] For a subset of individuals, Infutor also provides estimated demographic information (e.g., age, gender, and socioeconomic information). We do not incorporate this information in our analysis for two reasons: first, complete demographic data is only available for about half of all individuals; second, the reliability of estimated demographic information is concerning—as the data lacks documentation, we should not discard the possibility that demographic information has been imputed based on addresses—and contains known biases [22]. Dates are listed as month and year. For example, the data for one individual might consist of two addresses: “1 Main Street, Everytown, USA, 10000 (January 2010); 2 Cornelia Street, New York, New York, USA (December 2017)” followed by an initial date of January 2008 and an end date of December 2017. We describe the steps taken to (1) identify and clean Infutor data records within the scope of MIGRATE, (2) process the Infutor data into matrices documenting yearly flows between address pairs, and (3) map these address-level matrices to the CBG-level matrices $E^{(t)}$ mentioned in the main text. Table M1 summarizes the raw and processed data.
Cleaning address histories
@@
Processing address histories into annual address-address migration matrices
@@
Processing address-address migration matrices into CBG-CBG migration matrices
Fig. M1: Flowchart detailing the process of mapping addresses to CBGs. We start with all Infutor addresses and classify them into five address types. We remove addresses which lie within US territories but not US states (around .26% of the addresses). Incomplete addresses—those that do not have a precise street address line—as well as PO boxes and rural routes are mapped probabilistically to CBGs according to their ZIP code. The vast majority (90.4%) of addresses are clean, complete street addresses which are sent to the Census geocoder, which is able to map 81.26% of these addresses to a Census block group. If the Census geocoder fails to provide a match, we reattempt matching using ESRI’s ArcGIS geocoder; this achieves a lower match rate of 58.24%, in part because the sample it is applied to is more difficult to parse. In total, we are able to map 99.21% of all addresses in the original Infutor data: 83.33% to a precise latitude and longitude, and 15.89% to a ZIP code.
We transform the address-level migration matrices $A^{(t)}$ into CBG-level matrices $E^{(t)}$ by mapping each address either to a precise CBG where possible, or to a distribution over CBGs when only a ZIP code is available. We use the 2010 Census block group boundaries for all addresses to maintain geographic consistency. Figure M1 details this process.
Our address mapping pipeline is divided in two parts: mapping addresses via geocoding, and mapping addresses with a ZIP code-to-CBG crosswalk. We initially attempt to use one of two state-of-the art geocoders to map addresses to latitudes and longitudes: the publicly-available Census Bureau geocoder [27] and ESRI’s ArcGIS geocoder [26]. We first attempt to geocode an address via the Census geocoder; if this algorithm does not find a match, we submit the address to ArcGIS. If neither algorithm finds a match, we map the address to Census Tracts intersecting its ZIP code, weighting probabilities based on the share of residential ZIP code addresses represented by the population of each Census Tract. Addresses representing rural routes and postal boxes are directly mapped via ZIP code, as well as incomplete addresses. We use the HUD ZIP-to-tract crosswalk, with the date closest available to the last seen date of an address to account for the fact that ZIP code definitions may vary and are not nested within Census geographies [60]. We then distribute the probability to each Census block group within the tracts, again weighting by their relative population shares. Success rates for each of these steps are high and detailed in Figure M1.
@@
Table M1 reports a yearly breakdown of relevant Infutor data statistics throughout this pipeline. The number of individuals with active records in the data represents the number of individuals with an interval of activity containing each year at the start of the data processing; the number of CBG moves corresponds to the sum of off-diagonal entries of each matrix $E^{(t+1)}$ and measures the “effective size” of each year’s data: how many movers we actually account for at the beginning of the estimation procedure.
Year (1) Active records (2) US Population (3) Processed CBG Moves
2010 269,228,776 309,327,143 15,180,982
2011 260,064,921 311,583,481 11,793,806
2012 259,451,204 313,877,662 12,544,199
2013 252,933,953 316,059,947 13,829,056
2014 256,453,630 318,386,329 13,071,070
2015 241,295,957 320,738,994 14,022,249
2016 242,298,262 323,071,755 13,605,804
2017 193,479,001 325,122,128 12,684,184
2018 177,629,604 326,838,199 13,505,890
Table M1: Yearly Breakdown of Data Summaries. Population and addresses accounted by the Infutor and Census datasets during the analysis period (2010-2019). (1) Individuals have active records in a given year if the year falls within their interval of activity in the dataset. The number of active individuals decreases in recent years, as Infutor records tend to appear when a move is reported. Comparison of these values to (2) the Census population shows that Infutor under-counts the population at every given year. (3) The number of CBG-to-CBG moves parsed in our dataset. Note that an individual with two addresses will have at most one move, some addresses do not constitute residences, some individuals may have a single address in the year thus not move, etc.
M.2 Processing Census Data
The U.S. Census Bureau runs multiple programs with the aim of counting and estimating demographic and socioeconomic population data. Each of these programs may release data at different spatiotemporal granularity and, even if geographies or timespans agree, the data might lack internal consistency. We use a series of steps to clean, select, and process Census datasets. We describe our procedures for reconciling Census geographies, selecting which Census datasets to use as constraints, and how we process the Census data to deal consistently with births, deaths, emigration, and immigration.
Reconciling Census geographies
@@
Selecting Census datasets
@@
Table M2: ACS Coefficients of Variation. Average coefficient of variation (in %) of the non-zero estimates in each dataset. CVs were computed as the ratio of the standard error (obtained by dividing the margin of error by $1.645$) by the estimate [64]. We did not use flows released in 2010 (as they would report moves from 2009).
@@
We would like to avoid matching to datasets that were too noisily estimated. To decide which ACS datasets to use in our harmonization process, we examine their level of sampling error. We quantify the sampling error by the coefficient of variation (CV): the ratio of the standard deviation to the estimate value. Table M2 reports the mean coefficient of variation in ACS estimates. In general, the number of non-movers has very low CVs (mean below 2% for counties and below 1% for states); population estimates have considerably lower CVs than flow estimates; state-to-state flows have relatively high CVs (mean around 45% every year), and county-to-county flows have extremely high CVs (mean almost 90%). We opt not to match to county-county flows due to the imprecision of raw flow estimates, although we do use them as a validation. In-migration rates have much tighter CVs than the corresponding flows (mean around 15% and 5% for counties and states, respectively), and we also use those as validations.
@@
Accounting for natural population increase and international migration
@@
M.3 Harmonizing Infutor and Census Data
@@
Harmonizing with CBG population data
@@
Harmonizing with yearly state non-movers data
@@
Harmonizing with yearly state-to-state flow data
@@
Harmonizing with yearly county population data
@@
[…]
Supplementary information Supplementary Information is available detailing additional methods, validations, and findings.
Declarations
[…]
Data availability MIGRATE is available upon request for non-profit research use at migrate.tech.cornell.edu.
[…]
Code availability Code to reproduce our research findings is available on GitHub: https://github.com/gsagostini/MIGRATE.
Appendix A Additional details on validations
Here we provide additional details on MIGRATE validations. §A.1 provides additional validation metrics that complement Figure 2; §A.2 presents additional details on our held-out data validations; §A.3 provides details on related work documenting biases of Infutor, which is the subject of our Figure 3; and §A.4 presents additional socioeconomic and demographic dimensions on which we assessed biases.
A.1 Additional validation metrics
Fig. S1: Detailed flow validations of MIGRATE. To complement the findings shown in Figure 2, we present validations with respect to area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We show (a-c) Pearson correlations of estimates and selected ACS flows as well as (d-f) how much MIGRATE reduces RMSE compared to Infutor.
Figure S1 reports validation statistics for subsets of the migration matrices: in particular, area stayers (i.e., the diagonal entries of the migration matrix), area movers (i.e. the off-diagonal entries of the migration matrix), and raw in-migration (i.e. the number of people arriving at an area). We report these metrics at both the state and county level; state-to-state flows were used in the estimation procedure, county-to-county flows were held out. For all datasets, MIGRATE estimates are highly correlated with Census estimates ($\rho$ above $0.957$ in all cases) and reduce RMSE relative to the Infutor data.
A.2 Validating the estimation procedure using held-out data